OpenStudy (anonymous):

How do I solve and graph this inequality? 12x + 4 ≥ 16 or 3x – 5 ≤ –14

OpenStudy (anonymous):

Solve it like an equation but remember that the inequality changes if you divide or multiply by a negative number.

OpenStudy (anonymous):

I am new to this, could you help me do this problem? If I solve it like an equation... where do I start?

OpenStudy (anonymous):

Ok, so the first inequality is \[12x \ge 16-4\] \[12x \ge \] \[x \ge \frac{ 12 }{ 12 }\] \[x \ge 1\]

OpenStudy (anonymous):

\[12x \ge 12\]

OpenStudy (anonymous):

Okay, got it... so then is the second one x \[x \le 6\] ?

OpenStudy (anonymous):

\[3x \le -14+5\] \[3x \le -9\] \[x \le \frac{ -9 }{ 3 }\] \[x \le -3\]

OpenStudy (anonymous):

Oh, okay, I think I see what I did wrong here... so then to graph this...?

OpenStudy (anonymous):

Just ask questions if you have any. Graphing it is actually really simple, do you know what a numberline is?

OpenStudy (anonymous):